OpenStudy (anonymous):

Rewrite as a sum or difference. cos(5t)sin(8t)

OpenStudy (gw2011):

I'm not absolutely sure about this one but here is something that I think might answer this problem. (cos 5t)(sin 8t) is the same as (sin 8t)(cos 5t) Here is an equation that involves the above: sin (A+B)=(sinA)(cosB)+(cosA)(sinB) (sinA)(cosB)=sin (A+B)-(cosA)(sinB) Therefore, from the above A=8t and B=5t (sin8t)(cos5t)=sin (8t+5t)-(cos8t)(sin5t)